OpenStudy (anonymous):

write a expotential function y=abx for a graph that includes (0,5)(2,45) please explain

OpenStudy (anonymous):

go ahead

OpenStudy (watchmath):

Plugin \((0,5)\) we have \(5=a\). Plug in \((2,45\) and \(a=5\) \(45=5\cdot b^2\) \(b^2=9\) \(b=\pm 3\) So \(y=5\cdot 3^x\) or \(y=5\cdot (-3)^x\)

OpenStudy (anonymous):

well i think that you best take \[b=3\]

OpenStudy (anonymous):

\[b>0, b\neq 1\]

OpenStudy (anonymous):

reason being then you don't have to worry about something like \[(-3)^{\frac{1}{2}}=\sqrt{-3}\]

OpenStudy (watchmath):

Yes, just take b=3 :)